Wed Jun 3 signments a Circles: Unit test Graph the circle x² + y2 + 16x - 16y + 124 = 0. y 8 6 4 2 2 2 4 6 8 -4 -6
the andwer please thank you
Answer #1

e 2 + y2 + 168 -16y +124 - 0 we have to convert it tot into standard form of equation (x-4) ² + (y-k) ² = 12 where Chik is thAdd 64 with each bracket and equalize by putting is rett side. right (x ² + 16x+64) + Cy²- 160 +64) + 124 = 64164 (x+8)² + (y12 2) 8 10 radius (2) 6 -4 -2 -izo 7 2 46 8 x -8